OpenStudy (maheshmeghwal9):

A circular wire loop of radius 'a' carries a total charge 'Q' distributed uniformly over its length. A small length dL of the wire is cut off. What is the electric field at the center due to the remaining wire?

OpenStudy (maheshmeghwal9):



OpenStudy (turingtest):

the field that will act will be that for which there is no side symmetrically opposite it. That is, the other side that was cut out

OpenStudy (turingtest):

\[dQ=\lambda dL\]where \(\lambda\) is the linear charge density of the wire. Coulombs law gives\[d\vec E=k{\lambda\over a^2}dL\hat a\]

OpenStudy (maheshmeghwal9):

U r right @TuringTest :) But the answer given in my book is different & that is \[\huge{\color{red}{\vec E = \frac{Q \space dL}{8 \pi^2 \epsilon_0a^3}.}}\]

OpenStudy (maheshmeghwal9):

How can i get this answer ? Please anyone help:)

OpenStudy (maheshmeghwal9):

@experimentX Please help:)

OpenStudy (experimentx):

doesn't look easy

OpenStudy (maheshmeghwal9):

ya:(

OpenStudy (maheshmeghwal9):

Am I Right ?\[\LARGE{\vec E=\frac{\frac{Q}{2 \pi a}\times dL}{4 \pi \epsilon_0 \space a^2}}= \frac{Q \space dL}{8 \pi^2 \epsilon_0a^3}.\]

OpenStudy (experimentx):

Electric field due to a line element is \[ E = { k \lambda dl \over r^2} \\ l = 2 \pi r \implies dl = 2 \pi dr\]

OpenStudy (experimentx):

seems so ...

OpenStudy (maheshmeghwal9):

I have only thought about that. I think I must confirm this.

OpenStudy (experimentx):

probably it's the easiest way to do it.

OpenStudy (experimentx):

\[ E = { k \lambda dl \over r^2} \\ \lambda = {q\over 2 \pi r}\]

OpenStudy (maheshmeghwal9):

yea so i m right:)

OpenStudy (experimentx):

i had been thinking the straight way when the back way was more straight forward

OpenStudy (maheshmeghwal9):

hmm......

OpenStudy (maheshmeghwal9):

May i close this question? i mean have we gt the solution?

OpenStudy (experimentx):

yeah ...

OpenStudy (maheshmeghwal9):

ok. thank u very much:)

OpenStudy (experimentx):

originally ... the E at center is zero. it means that the E due to rest must have been balanced by dl ... so E due to dl should be equal to E due to rest.

OpenStudy (maheshmeghwal9):

ya .

OpenStudy (maheshmeghwal9):

@experimentX i have a confusion. would u plz help me ?

OpenStudy (experimentx):

where do you have?

OpenStudy (maheshmeghwal9):

I was saying that if i wanna have charge of remaining wire then i can subtract dL part's charge from net charge Q. Then can i get electric field due to the remaining wire?

OpenStudy (maheshmeghwal9):

@experimentX ?

OpenStudy (experimentx):

where did my answer go? dL will be equal to remaining part.

OpenStudy (maheshmeghwal9):

but i m asking that can't i get this way?

OpenStudy (experimentx):

yes you can ..

OpenStudy (maheshmeghwal9):

but actually i m nt getting would u plz show me?

OpenStudy (experimentx):

\[ Q - {Q \over 2 \pi a} dL\]

OpenStudy (maheshmeghwal9):

ya I did this but this is wrong:(\[\LARGE{\vec E=\frac{Q-\frac{Q}{2 \pi a}\times dL}{4 \pi \epsilon_0 \space a^2}}= \frac{2 \pi a Q -Q \space dL}{8 \pi^2 \epsilon_0a^3}.\]

OpenStudy (maheshmeghwal9):

It must be \[\huge{\color{red}{\vec E = \frac{Q \space dL}{8 \pi^2 \epsilon_0a^3}.}}\] Please help:)

OpenStudy (maheshmeghwal9):

I have problem in numerator.

OpenStudy (experimentx):

what are you trying to do? are you trying to get it the other way?

OpenStudy (maheshmeghwal9):

ya.

OpenStudy (maheshmeghwal9):

I have gt the remaining wire's charge so that i can get electric field due to that wire. but i m getting wrong:(